OpenStudy (anonymous):

-7/12 + y = 1/6 please show the stepsi know that the answer is 3/4 i just need a help to understand how , please and thank you:D

OpenStudy (anonymous):

adding 7/12 both sides y=7/12+1/6 now add y=(7*6+1*12)/12*6=3/4

OpenStudy (anonymous):

thank you :D